### v1

@@ 3.2 Re-parametrization
$\mathsf{F}\colon\mathbb{R}^{q}\times\mathbb{S}^{2}\times\mathbb{R}^{n\times 3}$ $\to\mathbb{R}^{n\times q}$
$(\mathbf{r},\mathbf{n},\mathsf{L})$ $\mapsto\mathsf{F}(\mathbf{r},\mathbf{n},\mathsf{L})=\mathsf{L}\,\mathbf{n}\,\mathbf{r}^{\top},$ (3)
@@
To ensure bijectivity of the mapping (2), we employ three distinct illumination vectors ($n=3$). Two additional constraints ensure bijectivity: (i) non-zero reflectance ($\|\mathbf{r}_{i,k}\|\neq 0$), valid as long as the surface is not perfectly black, and (ii) linear independence of illumination directions (non-singular $\mathsf{L}_{i,k}$). Under these conditions, the matrix $\mathsf{L}_{i,k}$ is invertible. If $q=1$, the inverse is given by $\mathbf{r}_{i,k}=\big{\|}\mathsf{L}_{i,k}^{-1}\mathbf{v}_{i,k}\big{\|},\quad\mathbf{n}_{i,k}=\frac{\mathsf{L}_{i,k}^{-1}\mathbf{v}_{i,k}}{\left\|\mathsf{L}_{i,k}^{-1}\,\mathbf{v}_{i,k}\right\|}$. If $q=3$, vectors $\mathbf{n}_{i,k}$ and $\mathbf{r}_{i,k}$ can be deduced from matrix $\mathsf{L}_{i,k}^{-1}\mathbf{v}_{i,k}=\mathbf{n}_{i,k}\mathbf{r}_{i,k}^{\top}$ using SVD.
@@ 3.3 Optimal Illumination Directions
Although the mapping $(\mathbf{r}_{i,k},\mathbf{n}_{i,k})\mapsto\mathbf{v}_{i,k}$ is bijective for any non-singular illumination matrix $\mathsf{L}_{i,k}$, not all choices are equally adapted to our task. Indeed, the input reflectances and normals being prone to estimation errors, radiance values can be considered “noisy”. Since we will next invert the PBR model from these noisy values, the conditioning of the illumination matrices will have a substantial impact on the accuracy of the numerical inversion.
We therefore chose the widely accepted configuration suggested by [20], which employs three illumination configurations with equal intensity, and directions spaced equally by $120^{\circ}$ in azimuth, each with a constant slant of $54.74^{\circ}$ relative to the normal vector $\mathbf{n}_{i,k}$. Alternative illumination arrangements, such as those embedded in SuperNormal [10], will be assessed in Section 5.3.
@@ 4 Surface Sweeping-based 3D Reconstruction
As a first proof of concept, let us now apply the re-parametrization we proposed in Section 3 to triangulation-based MVS, a method that estimates the depth for each pixel in a reference view by maximizing the photometric consistency in other (control) views. In particular, we will show that by leveraging the information provided by the normals, plane sweeping-based algorithms [16] can be turned into an exact surface sweeping-based method for integrating multi-view normal data, eliminating the inherent approximation errors due to local planar surface assumptions, e.g., fronto-parallel or slanted patches [25, 6, 73].
@@ 4.1 Objective Function
Our aim is to compute depth values for all pixels in the first view, considered as reference. Without loss of generality, this reference camera frame aligns with the world coordinate system. The $N-1$ other views then serve as control views for enforcing multi-view consistency on the reflectance and normals. The optimization of multi-view consistency is usually carried out locally, by splitting the pixels of the reference view into patches ${\mathcal{P}}\subset\mathbb{R}^{2}$ containing $m_{\mathcal{P}}$ pixels. We will denote by $\{\mathbf{v}_{1,j}\}_{j\in\{1,\dots,m_{P}\}}$ the re-parametrized reflectance and normals over this patch. Formulated in terms of this re-parametrization, finding the depth of the patch centre that maximizes multi-view consistency of reflectance and normal maps amounts to solving:
@@
where $\mathbf{v}_{i,j}(z)\in\mathbb{R}^{3\times q}$ stands for the re-parametrization in the $i$-th view, sampled using bilinear interpolation at a pixel position $\mathbf{p}_{i,j}(z)\in\mathbb{R}^{2}$ calculated given the depth hypothesis $z$. This position is obtained by first back-projecting the reference patch pixel $\mathbf{p}_{1,j}\in\mathcal{P}$ into 3D, and then projecting this 3D point into the control view:
@@
To proceed with the actual optimization of (4), the function $\pi_{z}^{-1}$ must now be made explicit, which typically involves a local geometric approximation (e.g., fronto-parallel or slanted patches). As we shall see next, the availability of normal information provides us with a natural way to avoid such an approximation.
@@ 4.2 Exact Multi-view Integration
  • Fronto-parallel patches (zero-order approximation): all the patch points are assumed to lie on a plane parallel to the image plane, located at the hypothesized depth $z$. The inverse projection is then simply $\pi_{z}^{-1}(\mathbf{p}_{1,j})=z\,\mathsf{K}^{-1}[\mathbf{p}_{1,j}^{\top}~{}1]^{\top}$, with $\mathsf{K}$ the intrinsics matrix of the reference camera (assumed known a priori). This is the simplest model, but it is often inaccurate for non-planar surfaces or slanted views [25, 26].
@@
By using this exact local surface representation, and the triangulation-based MVS framework (4) leveraging our re-parametrization, we not only enforce consistency in both appearance (reflectance) and detailed geometry (integrated normals) across views, but also avoid low-order approximation errors inherent to plane sweeping-based methods. This is empirically validated in the experiments presented hereafter.
@@ 4.3 Empirical Validation
To evaluate both the re-parametrization loss and the surface sweeping strategy, we generated a synthetic benchmark and compared our results against those obtained using traditional methods.
@@
Single-objective validation is first considered, by comparing our results against a more straightforward solution combining a reflectance discrepancy loss $\mathcal{L}_{\text{photo}}$ and a geometric one $\mathcal{L}_{\text{geom}}$, through a weighting hyperparameter $\mu$:
@@
Therein, tuning $\mu$ is often difficult, since its optimal value depends on the noise level. By unifying geometric and photometric cues into a single, parameter-free objective, the proposed re-parametrization circumvents this issue.
Figure 4 compares the mean depth error obtained using both approaches, in the presence of a fixed amount of additive Gaussian noise on the reflectance values (standard deviation: $1\%$ of the maximum reflectance value) and an increasing one on the normals. To focus the evaluation on the re-parametrization, the same slanted patch approximation was adopted in both cases. While the optimal value of $\mu$ depends on the noise level, our hyperparameter-free approach always represents a reasonable compromise in performance, regardless of the noise level.
@@
Surface sweeping validation is then be assessed, by comparing the various patch geometry approximations. To focus only on the sweeping method, this time we considered only our single-objective loss, and noise was added only to normals, not reflectance. To also compare patch-based MVS methods against more modern frameworks, we included in the comparison our previous implementation of RNb-NeuS [7], based on volumetric rendering. The results are presented in Figure 5.
@@
In the ideal, noiseless scenario, our proposed surface sweeping method leveraging normal integration achieves an essentially exact 3D reconstruction, demonstrating its capacity to eliminate the errors due to low-order patch geometry approximation. Such errors are indeed clearly visible on the first-order approximation (purple curve), and even more on the zero-order one (dark blue curve). The volumetric rendering approach (orange curve), on the other hand, presents a small, but non-null error, potentially stemming from inherent approximations within its neural implicit representation and volume rendering process which may include local planarity assumptions [82, 101].
Adding noise to the normals reveals critical differences in robustness. The accuracy of all patch-based variants degrades as noise increases. Most notably, the surface sweeping approach, while perfect initially, rapidly deteriorates to the same level as lower-order approximations. This indicates that the accuracy of the local surface derived from normal integration is highly vulnerable to inaccuracies in the input normal field. On the other hand, volumetric rendering demonstrates considerably better robustness to high noise levels. This robustness likely benefits from its global optimization framework and the implicit regularization provided by the optimization of an SDF.
Overall, these empirical findings validate our hyperparameter-free re-parametrization strategy, and the feasibility of exact 3D reconstruction in ideal conditions. However, given that state-of-the-art photometric stereo methods often yield normals with mean angular errors in the range $[5^{\circ},10^{\circ}]$ [31], we believe that robustness to noise prevails over theoretical accuracy. Therefore, in the next section we turn our attention to coupling our re-parametrization with volumetric rendering frameworks.
@@ 5 Volume Rendering-based 3D Reconstruction
In contrast with the previous section where we focused on a single depth map, the present one introduces a method for estimating a full 3D model (geometry and reflectance) that is consistent with the input normal and reflectance data. To do so, we embed the homogeneous radiance-based re-parametrization introduced in Section 3 into a unified objective function derived from neural volume rendering (NVR) principles. The actual implementation builds upon the NeuS2 framework [85], yielding a highly effective method for 3D reconstruction from multi-view normal and reflectance maps.
5.1 Surface Parametrization
We aim to infer a 3D model defined by two functions: a geometric map $f:\mathbb{R}^{3}\to\mathbb{R}$ and a photometric map $\boldsymbol{\rho}:\mathbb{R}^{3}\to\mathbb{R}^{q}$. Function $f$ represents the signed distance to the surface, such that the surface $\mathcal{S}$ is its zero-level set: $\mathcal{S}=\{\mathbf{x}\in\mathbb{R}^{3}\,|\,f(\mathbf{x})=0\}$. Function $\boldsymbol{\rho}$ assigns a reflectance value to each 3D point. Since the actual form of $\boldsymbol{\rho}$ must be consistent with the assumptions from Section 3, we will limit ourselves to the Lambertian model ($q=1$, or $q=3$ for RGB data), yet extensions to more complex BRDFs are conceivable as long as they are consistent with the PBR model used for re-parametrization.
@@ 5.2 Objective Function
For clarity within this subsection, we focus on data associated with a single camera view and omit the index $i$. Input reflectance and normal values are therefore denoted simply as $\{\mathbf{r}_{k}\}_{k}$ and $\{\mathbf{n}_{k}\}_{k}$, and their joint re-parametrization as $\{\mathbf{v}_{k}\}_{k}$.
The core idea of our NVR approach is to optimize the scene representation ($\boldsymbol{\rho},f$) such that its rendering, under the same conditions as in Section 3, matches the input-derived radiance vectors $\mathbf{v}_{k}$, by minimizing:
$\mathcal{L}_{\text{NVR}}^{p}(\boldsymbol{\rho},f)=\sum_{k=1}^{m}\|\tilde{\mathbf{v}}_{k}(\boldsymbol{\rho},f)-\mathbf{v}_{k}\|_{p}^{p},$ (8)
with $p\in\{1,2\}$ (this choice will be discussed in the experiments), and where $\tilde{\mathbf{v}}_{k}(\boldsymbol{\rho},f)$ is the NVR-based radiance at pixel $k$.
@@
$\tilde{\mathbf{v}}_{k}(\boldsymbol{\rho},f)=\int^{t_{1}}_{t_{0}}\!\!\!w(t,f(\mathbf{x}_{k}(t)))\,\mathbf{c}(\mathbf{x}_{k}(t),\boldsymbol{\rho},f)\,\mathrm{d}t,$ (9)
with $[t_{0},t_{1}]$ the integration range, and $w$ an occlusion-aware weighting function ensuring concentration around the surface [82].
Unlike the original NeuS which directly optimized apparent colour, we optimize the underlying surface properties by taking into account the PBR model $\mathsf{F}$ (it is crucial to use the same model as in (2)). Assimilating the unit outward surface normal to the gradient of the SDF, the apparent colour in (9) writes as:
$\mathbf{c}(\mathbf{x}_{k}(t),\boldsymbol{\rho},f)\!=\!\mathsf{F}\left(\boldsymbol{\rho}(\mathbf{x}_{k}(t)),\nabla f(\mathbf{x}_{k}(t)),\mathsf{L}_{k}\right),$ (10)
provided that $\|\nabla f\|=1$. To simplify the optimization process, we followed NeuS and relaxed this hard constraint into an eikonal regularizer (encouraging $\|\nabla f\|$ to be “close” to unity), controlled by some hyperparameter $\lambda\geq 0$:
@@
Combining (9) and (10) yields the NVR loss (8) for a single view. By averaging the contributions of all views, and adding the regularization (11) as well as the same silhouette consistency regularization as in NeuS [82], we obtain our complete loss function. This formulation enables end-to-end optimization of the SDF $f$ and reflectance $\boldsymbol{\rho}$ using gradient-based methods, typically by representing $f$ and $\boldsymbol{\rho}$ as MLPs and employing hierarchical sampling along rays, similar to NeuS [82].
@@ 5.3 SuperNormal: A Specific Case
Indeed, using (2) and (3) with $\mathbf{r}_{i,k}=[1]$ and $\mathsf{L}_{i,k}=\mathbf{I}_{3}$, the re-parametrized inputs simplify to the input normals: $\mathbf{v}_{i,k}=\mathbf{n}_{i,k}$ $\forall(i,k)$. Similarly, using (9) and (10) with $\boldsymbol{\rho}=\mathds{1}$ and $\mathsf{L}_{k}=\mathbf{I}_{3}$ yields the volumetric rendered normals: $\tilde{\mathbf{n}}_{k}(f):=\int^{t_{1}}_{t_{0}}w(t,f(\mathbf{x}_{k}(t)))\,\nabla f(\mathbf{x}_{k}(t))\,\mathrm{d}t$. With these two simplifications and $p=2$, the NVR loss (8) becomes:
@@
which is precisely the normal consistency loss introduced in SuperNormal. Upcoming experiments, presented in Section 6, will examine the performance obtained using this particular setting, in comparison with the proposed joint optimization over reflectance and normals using optimal triplets of light sources.

### v3

@@ 3.2 Re-parametrisation
$\mathsf{F}\colon\mathbb{S}^{2}\times\mathbb{R}^{q}\times\mathbb{R}^{n\times 3}$ $\to\mathbb{R}^{n\times q}$
$(\mathbf{n},\mathbf{r},\mathsf{L})$ $\mapsto\mathsf{F}(\mathbf{n},\mathbf{r},\mathsf{L})=\mathsf{L}\,\mathbf{n}\,\mathbf{r}^{\top},$ (3)
@@ 3.3 Optimal Illumination Directions
For this reason, we chose the widely accepted configuration suggested by [21], which employs three illumination configurations with equal intensity, and directions spaced equally by $120^{\circ}$ in azimuth, each with a constant slant of $54.74^{\circ}$ relative to the normal vector $\mathbf{n}_{i,k}$. Alternative illumination arrangements, such as those embedded in SuperNormal [10], will be assessed in Section 5.3.
@@ 4 Surface Sweeping-based 3D Reconstruction
As a first proof of concept, let us now apply the re-parametrisation we proposed in Section 3 to triangulation-based MVS, a method that estimates the depth for each pixel in a reference view by maximizing the photometric consistency in other (control) views. In particular, we will show that by leveraging the information provided by the normals, plane sweeping-based algorithms [17] can be turned into an exact surface sweeping-based method for integrating multi-view normal data, eliminating the inherent approximation errors due to local planar surface assumptions, e.g., fronto-parallel or slanted patches [26, 6, 75].
@@ 4.1 Objective Function
Our aim is to compute depth values for all pixels in the first view, considered as reference. Without loss of generality, this reference camera frame aligns with the world coordinate system. The N1N-1 other views then serve as control views for enforcing multi-view consistency on the normals and reflectance. The optimisation of multi-view consistency is usually carried out locally, by splitting the pixels of the reference view into patches ${\mathcal{P}}\subset\mathbb{R}^{2}$ containing $m_{\mathcal{P}}$ pixels. We will denote by $\{\mathbf{v}_{1,j}\}_{j\in\{1,\dots,m_{P}\}}$ the re-parametrised normals and reflectance over this patch. Formulated in terms of this re-parametrisation, finding the depth of the patch centre that maximises multi-view consistency of normal and reflectance maps amounts to solving:
@@
where $\mathbf{v}_{i,j}(z)\in\mathbb{R}^{3\times q}$ stands for the re-parametrisation in the $i$-th view, sampled using bilinear interpolation at a pixel position $\mathbf{p}_{i,j}(z)\in\mathbb{R}^{2}$ calculated given the depth hypothesis $z$. This position is obtained by first back-projecting the reference patch pixel $\mathbf{p}_{1,j}\in\mathcal{P}$ into 3D, and then projecting this 3D point into the control view:
@@
To proceed with the actual optimisation of (4), the function $\pi_{z}^{-1}$ must now be made explicit, which typically involves a local geometric approximation (e.g., fronto-parallel or slanted patches). As we shall see next, the availability of normal information provides us with a natural way to avoid such an approximation.
@@ 4.2 Exact Multi-view Integration
  • Fronto-parallel patches (zero-order approximation): all the patch points are assumed to lie on a plane parallel to the image plane, located at the hypothesised depth $z$. The inverse projection is then simply $\pi_{z}^{-1}(\mathbf{p}_{1,j})=z\,\mathsf{K}^{-1}[\mathbf{p}_{1,j}^{\top}~1]^{\top}$, with $\mathsf{K}$ the intrinsics matrix of the reference camera (assumed known a priori). This is the simplest model, but it is often inaccurate for non-planar surfaces or slanted views [26, 27].
@@
By using this exact local surface representation, and the triangulation-based MVS framework (4) leveraging our re-parametrisation, we not only enforce consistency in both detailed geometry (integrated normals) and appearance (reflectance) across views, but also avoid low-order approximation errors inherent to plane sweeping-based methods. This is empirically validated in the experiments presented hereafter.
@@ 4.3 Empirical Validation
To evaluate both the re-parametrisation loss and the surface sweeping strategy, we generated a synthetic benchmark and compared our results against those obtained using traditional methods.
@@
Single-objective validation is first considered, by comparing our results against a more straightforward solution combining a geometric loss $\mathcal{L}_{\text{geom}}$ and a reflectance discrepancy one $\mathcal{L}_{\text{photo}}$, through a weighting hyperparameter $\mu$:
@@
Therein, tuning $\mu$ is often difficult, since its optimal value depends on the noise level. By unifying geometric and photometric cues into a single, parameter-free objective, the proposed re-parametrisation circumvents this issue.
Figure 4 compares the mean depth error obtained using both approaches, in the presence of an increasing amount of additive Gaussian noise on the orientation of the normals and a fixed one on the reflectance values (standard deviation: $1\%$ of the maximum reflectance value). To focus the evaluation on the re-parametrisation, the same slanted patch approximation was adopted in both cases. While the optimal value of $\mu$ depends on the noise level, our hyperparameter-free approach always represents a reasonable compromise in performance, regardless of the noise level.
@@
To focus solely on the evaluation and comparison of sweeping methods, this time we considered only our single-objective loss, and noise was added only to normals, not to reflectance. To also compare patch-based MVS methods against more modern frameworks, we included in the comparison our previous implementation of RNb-NeuS [7], based on volumetric rendering. The results are presented in Figure 5.
@@
In the ideal, noiseless scenario, our proposed surface sweeping method leveraging normal integration achieves an essentially exact 3D reconstruction, demonstrating its capacity to eliminate the errors due to low-order patch geometry approximation. Such errors are indeed clearly visible on the first-order approximation (purple curve), and even more on the zero-order one (dark blue curve).
Adding noise to the normals reveals critical differences in robustness. The accuracy of all patch-based variants degrades as noise increases. Most notably, the surface sweeping approach, while perfect initially, rapidly deteriorates to the same level as lower-order approximations. This indicates that the accuracy of the local surface derived from normal integration is highly vulnerable to inaccuracies in the input normal field. On the other hand, volumetric rendering demonstrates considerably better robustness to high noise levels. This robustness likely benefits from its global optimisation framework and the implicit regularisation provided by the optimisation of an SDF.
Overall, these empirical findings validate our hyperparameter-free re-parametrisation strategy, and the feasibility of exact 3D reconstruction in ideal conditions. However, given that state-of-the-art photometric stereo methods often yield normals with mean angular errors in the range $5^{\circ},10^{\circ}] [$32], we believe that robustness to noise prevails over theoretical accuracy. Therefore, in the next section we turn our attention to coupling our re-parametrisation with volumetric rendering frameworks.
@@ 5 Volume Rendering-based 3D Reconstruction
In contrast with the previous section where we focused on a single depth map, the present one introduces a method for estimating a full 3D model (geometry and reflectance) that is consistent with the input normal and reflectance data. To do so, we embed the homogeneous radiance-based re-parametrisation introduced in Section 3 into a unified objective function derived from neural volume rendering (NVR) principles. The actual implementation builds upon the NeuS2 framework [88], yielding a highly effective method for 3D reconstruction from multi-view normal and reflectance maps.
5.1 Surface parametrisation
We aim to infer a 3D model defined by two functions: a geometric map $f:\mathbb{R}^{3}\to\mathbb{R}$ and a photometric map $\boldsymbol{\rho}:\mathbb{R}^{3}\to\mathbb{R}^{q}$. Function $f$ represents the signed distance to the surface, such that surface $\mathcal{S}$ is its zero-level set: $\mathcal{S}=\{\mathbf{x}\in\mathbb{R}^{3}\,|\,f(\mathbf{x})=0\}$. Function $\boldsymbol{\rho}$ assigns a reflectance value to each 3D point. Since the actual form of $\boldsymbol{\rho}$ must be consistent with the assumptions from Section 3, we will limit ourselves to the Lambertian model ($q=1$, or $q=3$ for RGB data), yet extensions to more complex BRDFs are conceivable as long as they are consistent with the PBR model used for re-parametrisation.
@@ 5.2 Objective Function
For clarity within this subsection, we focus on data associated with a single camera view and omit the index $i$. Input normal and reflectance values are therefore denoted simply as $\{\mathbf{n}_{k}\}_{k}$ and $\{\mathbf{r}_{k}\}_{k}$, and their joint re-parametrisation as $\{\mathbf{v}_{k}\}_{k}$.
The core idea of our NVR approach is to optimise the scene representation ($f,\boldsymbol{\rho}$) such that its rendering, under the same conditions as in Section 3, matches the input-derived radiance vectors $\mathbf{v}_{k}$, by minimising:
$\mathcal{L}_{\text{NVR}}^{p}(f,\boldsymbol{\rho})=\sum_{k=1}^{m}\|\tilde{\mathbf{v}}_{k}(f,\boldsymbol{\rho})-\mathbf{v}_{k}\|_{p}^{p},$ (8)
with $p\in\{1,2\}$ (this choice will be discussed in the experiments), and where $\tilde{\mathbf{v}}_{k}(f,\boldsymbol{\rho})$ is the NVR-based radiance at pixel $k$.
@@
$\tilde{\mathbf{v}}_{k}(f,\boldsymbol{\rho})=\int^{t_{1}}_{t_{0}}\!\!\!w(t,f(\mathbf{x}_{k}(t)))\,\mathbf{c}(\mathbf{x}_{k}(t),f,\boldsymbol{\rho})\,\mathrm{d}t,$ (9)
with $t_{0},t_{1}]$ the integration range, and $w$ an occlusion-aware weighting function ensuring concentration around the surface [85].
Unlike the original NeuS which directly optimised apparent colour, we optimise the underlying surface properties by taking into account the PBR model $\mathsf{F}$ (it is crucial to use the same model as in (2)). Assimilating the unit outward surface normal to the gradient of the SDF, the apparent colour in (9) writes as:
$\mathbf{c}(\mathbf{x}_{k}(t),f,\boldsymbol{\rho})\!=\!\mathsf{F}\left(\nabla f(\mathbf{x}_{k}(t)),\boldsymbol{\rho}(\mathbf{x}_{k}(t)),\mathsf{L}_{k}\right),$ (10)
provided that $\|\nabla f\|=1$. To simplify the optimisation process, we followed NeuS and relaxed this hard constraint into an eikonal regulariser (encouraging $\|\nabla f\|$ to be “close” to unity), controlled by some hyperparameter $\lambda\geq 0$:
@@
Combining (9) and (10) yields the NVR loss (8) for a single view. By averaging the contributions of all views, and adding the regularisation (11) as well as the same silhouette consistency regularisation as in NeuS [85], we obtain our complete loss function. This formulation enables end-to-end optimisation of the SDF $f$ and reflectance $\boldsymbol{\rho}$ using gradient-based methods, typically by representing $f$ and $\boldsymbol{\rho}$ as MLPs and employing hierarchical sampling along rays, similar to NeuS [85].
@@ 5.3 SuperNormal: A Specific Case
Indeed, using (2) and (3) with $\mathbf{r}_{i,k}=[1]$ and $\mathsf{L}_{i,k}=\mathbf{I}_{3}$, the re-parametrised inputs simplify to the input normals: $\mathbf{v}_{i,k}=\mathbf{n}_{i,k}$ $\forall(i,k)$. Similarly, using (9) and (10) with $\boldsymbol{\rho}=\mathds{1}$ and $\mathsf{L}_{k}=\mathbf{I}_{3}$ yields the volumetric rendered normals: $\tilde{\mathbf{n}}_{k}(f):=\int^{t_{1}}_{t_{0}}w(t,f(\mathbf{x}_{k}(t)))\,\nabla f(\mathbf{x}_{k}(t))\,\mathrm{d}t$. With these two simplifications and $p=2$, the NVR loss (8) becomes:
@@
which is precisely the normal consistency loss introduced in SuperNormal. Upcoming experiments, presented in Section 6, will examine the performance obtained using this particular setting, in comparison with the proposed joint optimisation over normals and reflectance using optimal triplets of light sources.